OpenStudy (anonymous):

Write the expression in standard form. 3/(3-12i)

OpenStudy (anonymous):

1/(1-4i)

ganeshie8 (ganeshie8):

\(\large \frac{3}{3-12i}\)

ganeshie8 (ganeshie8):

is it like above ? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

the answer choices are -1/17 + 4/17i , 1/17-4/17i, -1/17-4/17i, and 1/17 + 4/17i

OpenStudy (jdoe0001):

write it in standard form, in short means, get rid of the complex expression at the bottom how? well, multiply top and bottom by the \(\bf \text{conjugate}\) of the bottom

OpenStudy (anonymous):

I don't understand how to do that when it has complex zeros aka the i

OpenStudy (anonymous):

I see what you explained but I don't understand the relationship of that and the answer choices

OpenStudy (anonymous):

Would it be 1/17 - 4/17i?

ganeshie8 (ganeshie8):

complex conjugate of \(3-12i\) is \(3+12i\) jdoe is asking u to multiply top and bottom wid this

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

\(\large \frac{3}{3-12i} \) \(\large \frac{3}{3-12i} \times \frac{3+12i}{3+12i} \)

OpenStudy (anonymous):

but how does multiplying the "i" work?

ganeshie8 (ganeshie8):

\(\large \frac{3}{3-12i} \) \(\large \frac{3}{3-12i} \times \frac{3+12i}{3+12i} \) \(\large \frac{3(3+12i)}{(3-12i)(3+12i)} \)

ganeshie8 (ganeshie8):

for now, i may use this :- \(\large (a+bi)(a-bi) = a^2+b^2\)

ganeshie8 (ganeshie8):

*u/we

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

use that and simplify denominator

ganeshie8 (ganeshie8):

\(\large \frac{3}{3-12i} \) \(\large \frac{3}{3-12i} \times \frac{3+12i}{3+12i} \) \(\large \frac{3(3+12i)}{(3-12i)(3+12i)} \) \(\large \frac{3(3+12i)}{3^2+12^2} \)

ganeshie8 (ganeshie8):

simplify completely

ganeshie8 (ganeshie8):

\(\large \frac{3}{3-12i} \) \(\large \frac{3}{3-12i} \times \frac{3+12i}{3+12i} \) \(\large \frac{3(3+12i)}{(3-12i)(3+12i)} \) \(\large \frac{3(3+12i)}{3^2+12^2} \) \(\large \frac{3 \times 3(1+4i)}{153} \)

OpenStudy (anonymous):

I got 1+2i/17 . It's supposed to be something else though.

ganeshie8 (ganeshie8):

\(\large \frac{3}{3-12i} \) \(\large \frac{3}{3-12i} \times \frac{3+12i}{3+12i} \) \(\large \frac{3(3+12i)}{(3-12i)(3+12i)} \) \(\large \frac{3(3+12i)}{3^2+12^2} \) \(\large \frac{3 \times 3(1+4i)}{153} \) \(\large \frac{9(1+4i)}{9 \times 17} \)

OpenStudy (anonymous):

so 1/17 + 4/17i would be correct?

ganeshie8 (ganeshie8):

\(\large \frac{3}{3-12i} \) \(\large \frac{3}{3-12i} \times \frac{3+12i}{3+12i} \) \(\large \frac{3(3+12i)}{(3-12i)(3+12i)} \) \(\large \frac{3(3+12i)}{3^2+12^2} \) \(\large \frac{3 \times 3(1+4i)}{153} \) \(\large \frac{\cancel{9}(1+4i)}{\cancel{9} \times 17} \) \(\large \frac{1+4i}{17} \)

OpenStudy (anonymous):

yes, thank you again!

ganeshie8 (ganeshie8):

so 1/17 + 4/17i would be correct? yup, wud be correct :)

ganeshie8 (ganeshie8):

np. you're wlcme !